# Free field construction of Heterotic string compactified on Calabi-Yau manifolds of Berglund-Hubsch type in the Batyrev-Borisov combinatorial approach

### v5

@@ 8 Massless right movers and E(6) gauge symmetry
This means that we have to choose and keep among the vertices that are BRST cohomology, $D_{\vec{m}}$ and $D_{\vec{n}}$, only those that are mutually local with the currents $E(8)$ and $E(6)$.
@@
9 Massless vertices in explicit form.
@@
In this section we explicitly represent complete vertices for massless physical states as products of suitable left and right vertices, where we omit the factor $\exp\left(\imath p_{\mu}x^{\mu}\right)$ in order to shorten notations.
@@ 9.1 The gravity supermultiplets
First, we obtain expressions for the vertices of the gravitational supermultiplet, including the graviton vertex
@@
9.2 The gauge supermultiplets
We also get the following expressions for the vertices of vector gauge supermultiplets. The currents of $E(8)$ are
@@ 9.2 The gauge supermultiplets
where the vectors $\vec{\epsilon}$ are the roots of $E(8)$ algebra;
@@
$J^{\pm}_{\omega}(\bar{z})=\exp{(\imath\omega_{\alpha}\Phi_{\alpha}\pm\frac{1}{2}\bar{H}_{CY})(\bar{z})}$ (67)
@@
$J^{\pm}_{\dot{\omega}}(\bar{z})=\exp{(\imath\dot{\omega}_{\alpha}\Phi_{\alpha}\pm\frac{1}{2}\bar{H}_{CY})(\bar{z})},$
@@
As a result we obtain the $32$ additional currents of $E(6)$ algebra
@@
that together with the other $46$ currents of $SO(10)$ and $U(1)$ form the $78$-dimensional adjoint representation of E(6) algebra.
9.3 $27$ and $\overline{27}$ supermultiplets
The third type are the vertex operators of $27$ supermultiplet of $E(6)$ algebra whose right-moving factors belong to the 1-dimensional, 10-dimensional and 16- dimensional representations of $S(10)$ algebra. The left-moving factors of all these vertex operators belong to spinor reppresentation of $N=1$ Super Poincare algebra.
Thus, the vertex operators of the $27$ supermultiplet include a 1-dimensional representation of $SO(10)$
$V_{\vec{\sigma},\vec{m}}(z,\bar{z})=\exp\left(-\frac{1}{2}\phi+\imath{\vec{\sigma}}\cdot\vec{\tilde{H}}+\frac{1}{2}H^{L}_{CY}\right)\times\exp{(\bar{H^{R}}_{CY}+\vec{m}\cdot\vec{X}^{-})},$ (69)
@@ 9.3 $27$ and $\overline{27}$ supermultiplets
$V_{\vec{\sigma},\vec{m}}^{\vec{\Lambda}}(z,\bar{z})=\exp\left(-\frac{1}{2}\phi+\imath{\vec{\sigma}}\cdot\vec{\tilde{H}}+\frac{1}{2}H^{L}_{CY}\right)\times\exp{(i\vec{\Lambda}\cdot\vec{\Phi}+\vec{m}\cdot\vec{X}^{-})},$ (70)
@@
$V_{\vec{\sigma},\vec{m}}^{\dot{\vec{\omega}}}(z,\bar{z})=\exp\left(-\frac{1}{2}\phi+\imath{\vec{\sigma}}\cdot\vec{\tilde{H}}+\frac{1}{2}H^{L}_{CY}\right)\times\exp{\left(i\vec{\dot{\omega}}\cdot\vec{\Phi}+\frac{1}{2}\bar{H^{R}}_{CY}+\vec{m}\cdot\vec{X}^{-}\right)},$ (71)
@@
9.4 Singlet supermultiplets
The final, fourth class of vertices that satisfies all the requirements of our construction is the $N=1$ supermultiplets, which are $E(8)\times E(6)$ singlets.
The first type of the singlet vertices is
$V_{\vec{\sigma},\vec{m},\vec{n}}(z,\bar{z})=\exp\left(-\frac{1}{2}\phi+\imath{\vec{\sigma}}\cdot\vec{\tilde{H}}+\frac{1}{2}H^{L}_{CY}\right)\times\exp(\vec{m}\cdot\vec{X}^{-}+\vec{n}\cdot\vec{X}^{+}),$ (72)
where the vectors $\vec{m}\in\Delta^{+}$, $\vec{n}\in\Delta^{-}$ and their product $\vec{m}\cdot\vec{n}=0$.
The second type of the singlet vertices is
$V_{\vec{\sigma},\vec{m}}(z,\bar{z})=\exp\left(-\frac{1}{2}\phi+\imath{\vec{\sigma}}\cdot\vec{\tilde{H}}+\frac{1}{2}H^{L}_{CY}\right)\times\sum_{i}m_{i}\exp{(-i\bar{H}_{i}+\vec{m}\cdot\vec{X}^{-})},$ (73)
where $\vec{m}\in\Delta^{+}$.
The third type of the singlet vertices is
$V_{\vec{\sigma},\vec{n}}(z,\bar{z})=\exp\left(-\frac{1}{2}\phi+\imath{\vec{\sigma}}\cdot\vec{\tilde{H}}+\frac{1}{2}H^{L}_{CY}\right)\times\sum_{i}n_{i}\exp{(i\bar{H}_{i}+\vec{n}\cdot\vec{X}^{+})},$ (74)
where $\vec{n}\in\Delta^{-}$.
@@ 10 Conclusion
In particular, the total number of singlets in the Quintic case is 326, which coincides with the number obtained by Gepner Gep .
In this paper we have not discussed the modular properties of the partition function. We expect, that similar to BEL2 the condition of mutual locality would imply modular invariance authomatically. But this question deserves further studies and will be considered elsewhere.

### v6

@@ 8 Massless right movers and E(6) gauge symmetry
This means that we have to choose and keep among the vertices that are BRST and $D_{\vec{1},\vec{1}}$ cohomology, only those that are mutually local with the currents $E(8)$ and $E(6)$.
@@
9 Complete massless vertices in explicit form
@@
In this section we explicitly represent complete vertices for massless physical states as products of suitable left and right vertices. We will omit the factor $\exp\left(\imath p_{\mu}x^{\mu}\right)$ in order to shorten notations.
In choosing suitable products of left and right vertices, we follow D. Gepner’s approach in Gep to achieve modular invariance in constructing heterotic string theory.
Namely, we mutiply the left movers of [0] by the right movers of [V] and vice versa, and also mutiply the left movers of [S] by the right movers of [C] and vice versa.
As a result, we obtain only subsets of complete vertices belonging to each of the 4 types of massless physical states. Other complete vertices, which are nothing more than partners of subset vertices, can be obtained by applying $N=1$ Super-Poincaré actions to the left and $E(8)\times(6)$ symmetries to the right vertices.
@@ 9.1 The gravity supermultiplets
The first class of the massless physical states includes the gravitational supermultiplet.
These vertices are the vector representation of Space-Time $N=1$ Super-Poincare and the singlet representation of gauge group $E(8)\times E(6)$.
@@
9.2 The supermultiplets for the gauge group $E(8)\times E(6)$
The second class includes the adjoint representations of the gauge groups $E(8)$ and $E(6)$.
The vertices of $E(8)$ are
@@ 9.2 The supermultiplets for the gauge group $E(8)\times E(6)$
where the vectors $\vec{\epsilon}$ are the roots of $E(8)$ algebra.
@@
$J^{+}_{\omega}(\bar{z})=\exp{(\imath\omega_{\alpha}\Phi_{\alpha}+\frac{1}{2}\bar{H}_{CY})(\bar{z})}$ (65)
@@
$J^{-}_{\dot{\omega}}(\bar{z})=\exp{(\imath\dot{\omega}_{\alpha}\Phi_{\alpha}-\frac{1}{2}\bar{H}_{CY})(\bar{z})},$
@@
As a result we obtain the $32$ additional vertices of $E(6)$ algebra
@@
that together with the other $46$ vertices of $SO(10)$ and $U(1)$ form the $78$-dimensional adjoint representation of E(6) algebra.
The second class also includes the superpartners of these vertices listed above.
9.3 The supermultiplets $27$ and $\overline{27}$ of group $E(6)$
The third class is the representation $27$ of algebra $E(6)$ . The set of its vertices includes a one-dimensional representation $SO(10)$
$V_{\vec{m}}(z,\bar{z})=\exp\left(-\phi+\vec{m}\cdot\vec{X}^{-}\right)\times\exp{(-\bar{H^{R}}_{CY}+\vec{m}\cdot\vec{X}^{-})},$ (67)
@@ 9.3 The supermultiplets $27$ and $\overline{27}$ of group $E(6)$
$V_{\vec{\Lambda},\vec{m}}(z,\bar{z})=\exp\left(-\phi+\vec{m}\cdot\vec{X}^{-}\right)\times\exp{(i\vec{\Lambda}\cdot\vec{\Phi}+\vec{m}\cdot\vec{X}^{-})},$ (68)
@@
$V_{\vec{\omega},\vec{m}}(z,\bar{z})=\exp\left(-\phi+\vec{m}\cdot\vec{X}^{-}\right)\times\exp{\left(i\vec{\omega}\cdot\vec{\Phi}-\frac{1}{2}\bar{H^{R}}_{CY}+\vec{m}\cdot\vec{X}^{-}\right)},$ (69)
@@
9.4 The singlet supermultiplets of $E(8)\times E(6)$
The fourth class the massless states is the vector representation of N=1 Super-Poincare algebra and the singlet representation of $E(8)\times E(6)$ algebra.
Its first set of these vertices is
$V_{\vec{\lambda},\vec{m},\vec{n}}(z,\bar{z})=\exp\left(-\phi+\imath{\vec{\lambda}}\cdot\vec{\tilde{H}}+\frac{1}{2}H^{L}_{CY}\right)\times\exp(\vec{m}\cdot\vec{X}^{-}+\vec{n}\cdot\vec{X}^{+}),$ (70)
where the vectors $\vec{\lambda}=(\pm 1,0),(0,\pm 1)$, $\vec{m}\in\Delta^{+}$, $\vec{n}\in\Delta^{-}$ and their product $\vec{m}\cdot\vec{n}=0$.
The second set is
$V_{\vec{\lambda},\vec{m}}(z,\bar{z})=\exp\left(-\phi+\imath{\vec{\lambda}}\cdot\vec{\tilde{H}}+\frac{1}{2}H^{L}_{CY}\right)\times\sum_{i}m_{i}\exp{(-i\bar{H}_{i}+\vec{m}\cdot\vec{X}^{-})},$ (71)
where the vectors $\vec{\lambda}=(\pm 1,0),(0,\pm 1)$ and $\vec{m}\in\Delta^{+}$.
The third set is
$V_{\vec{\lambda},\vec{n}}(z,\bar{z})=\exp\left(-\phi+\imath{\vec{\lambda}}\cdot\vec{\tilde{H}}+\frac{1}{2}H^{L}_{CY}\right)\times\sum_{i}n_{i}\exp{(i\bar{H}_{i}+\vec{n}\cdot\vec{X}^{+})},$ (72)
where $\vec{\lambda}=(\pm 1,0),(0,\pm 1)$ and $\vec{n}\in\Delta^{-}$.
@@ 10 Conclusion
In particular, the total number of singlets in the Quintic case is 326, which coincides with the number obtained in Gep .
In choosing suitable products of left and right vertices, we follow D. Gepner approach Gep to achieve the modular invariance in constructing heterotic string theory. Although the correctness of this approach has been proved for the CY compactification used in Gep , it would be interesting to prove the fulfillment of modular symmetry in the wider class of CY considered in this paper.